You made a solution of sodium benzoate by dissolving 1.9 grams of sodium benzoate in 20 mL of distilled water. Calculate its molarity. (MW of sodium benzoate is 144 g)
Answer #1

We have

Mass of sodium benzoate = 1.9 grams

Molar mass of sodium benzoate = 144 g/mol

Volume of distilled water = 20 mL = 20/1000 L = 0.02 L

We know the formula of mole

Mole = mass in g / Molar mass

So lets calculate moles of sodium benzoate

Moles (n) of sodium benzoate = mass of sodium benzoate (m)/ Molar mass of sodium benzoate (MW)

= 1.9 g / 144 g/mol = 0.0132 mol

Now, We know the formula of Molarity

Molarity (M) = number of moles (n) / Volume in Liter (V)

So

Molarity of sodium benzoate = moles of sodium benzoate / Volume of Solution

= 0.0132 mol / 0.02 L = 0.66 M

So our answer = 0.66 M